OpenStudy (anonymous):

The equation of a line is shown below: y = - x - 7 What is the equation of a line which is parallel to this line and passes through (1, 79)?

OpenStudy (anonymous):

y = -x + 80 y = x - 80 y = 7x - 80 y = -7x + 80

OpenStudy (anonymous):

@stoopidfaka